### v1

@@ 3.4 Early Aborting Robotic Picking upon Grasp Adjustment and Slip Prediction
Empty grasp was a common failure mode in robotic strawberry harvesting during the deflating and snap-off stages. During the deflating stage, empty grasps occurred when the strawberry was not successfully enclosed by the end-effector, primarily due to localization errors or unexpected fruit motion. In the snap-off stage, strawberry slippage occurred when the picked strawberry unintentionally slipped, often due to a non-optimal end-effector pose or inadequate gripping force. To diagnose such faults effectively, a simple yet efficient method involved the installation of a miniature camera (JTS302, JiuTan, China) with a 3 mm diameter at the bottom of the end-effector, as shown in Fig. 6. The micro-optical camera directly captured real-time images of the gripper area, enabling visual monitoring of strawberry presence.
@@
Once the deflating stage began, a MobileNet V3-Small was used to determine whether a strawberry was present in the gripper, as shown in Fig. 7. If a strawberry was detected, the robotic arm continued executing the subsequent picking actions. Otherwise, the robot skipped the deflating and placing stages of the picking sequence. Instead, it immediately selected a new target strawberry and proceeded to the next picking cycle. The process flow, indicated by the blue line, was illustrated in Fig. 3.
@@
For the slippage scenario, a time-series LSTM classifier was developed to predict whether the strawberry would slip from the end-effector during the snap-off stage. The LSTM classifier architecture was summarized in Table 1. The input layer received a feature vector of dimension 5 from SRR-Net, which was followed by a hidden LSTM layer with 32 units that captured the temporal dependencies in the sequential data. The hidden representation was then mapped through a linear layer of size 32 and finally projected to a single output neuron that produced the classification result.
@@
input 5
hidden 32
linear 32
output 1
The input of the LSTM classifier consisted of a sequence of features from the past 10 frames, represented as <<normalized strawberry area, normalized gripper area, normalized background area, $w$, $h$ >>. The normalized strawberry area, gripper area, and background area referred to the proportion of pixels in the image corresponding to the strawberry, the gripper, and the background, respectively. $w$ and $h$ represented the width and height of the strawberry. The model outputted the probability of slippage over the following three frames. If the probability fell below the minimum threshold, the strawberry remained attached to the end-effector after the snap-off motion, indicating a successful pick. Conversely, if the probability was equal to or exceeded the maximum threshold, the strawberry had slipped, indicating a failure in the snap-off stage. When the probability fell between the minimum and maximum thresholds, the strawberry was in the process of slipping from the gripper.
@@
To enhance reliability, a time-stability rule was implemented. If slippage was predicted in three consecutive instances, an early abort signal was triggered, the placing stage was skipped, and the next picking cycle began. If the number of slipping instances was greater than zero but less than three, a secondary picking mechanism was activated, bypassing the inflating and approaching and swallowing stages, with the gripper inflating and deflating to re-grasp the strawberry during the snap-off actions. When no slip was predicted, the robot proceeded with the remaining steps until the cycle was completed.
@@ 3.5 Paradigm of HarvestFlex Visual Fault Diagnosis and Self-Recovery
The relative positional error compensation method, incorporating visual feedback SRR-Net, and the early abort strategy based on MobileNet V3-Small and the LSTM-based strawberry slippage prediction were implemented as follows. The gripper then inflated and approached a location directly beneath the target strawberry picking point. At this position, the relative error between the gripper and the actual strawberry position was calculated. If the error exceeded a predefined tolerance threshold, a compensated picking point was generated to correct the deviation before proceeding.
Once aligned, the gripper executed swallowing and deflating motions to enclose the strawberry. During this process, a micro-optical camera mounted on the gripper captured images, which were used to determine whether the strawberry was successfully gripped based on MobileNet V3-Small classifier. If no strawberry was detected in the gripper for three consecutive frames, the robot executed descending and homing actions to initiate a new picking cycle. Conversely, it continuously performed the bending action while running a time-series LSTM classifier to predict strawberry slippage.
To improve prediction stability, a time-stability rule was applied. When the probability exceeded the maximum threshold for three consecutive instances, the robot aborted the snap-off stage, skipped the placing stage, and returned to its homing position, as illustrated by the red lines in Fig. 9. When the probability remained below the minimum threshold, the robot continued the snap-off motion before proceeding to the next stage, as illustrated by the black line in Fig. 9. Meanwhile, if the probability fell between the maximum and minimum thresholds for one or two instances, a new picking workflow was initiated, bypassing the approaching and swallowing stages. In this case, the gripper was inflated and deflated to re-grasp the strawberry during the deflating and snap-off actions, followed by the descending, placing, and homing steps, as illustrated by the orange line in Fig. 9.
@@ 4.1 Evaluation Metrics
For the relative error between the end-effector and the picked strawberry instance, the visual relative errors along the $x$-, $y$-, and $z$- axes under the robotic arm coordinate system were denoted as $\Delta x$, $\Delta y$, and $\Delta z$, respectively.
Due to the bending motion required during the snap-off stage, the end-effector’s ascent height could exceed that of the target point. Therefore, the error calculation was simplified from the $x$-, $y$-, and $z$-axes to only the $x$-, and $y$- axes. Similarly, the corresponding physical errors along the $x$-, and $y$-axes were denoted as $\Delta x_{w}$ and $\Delta y_{w}$. The physical error after compensation, denoted as $E_{x}$ and $E_{y}$, was measured to validate the effectiveness of the relative error compensation method. Finally, mean value $\bar{e}$ was used to evaluate the accuracy of the relative error and physical error before compensation and for each coordinate axis and can be formulated as:
@@ 4.2 Implementation Details
To evaluate the effectiveness of the proposed method, experiments were conducted on the FaultData, GraspData, and SnapData benchmarks, with GraspData and SnapData constructed using artificial strawberries. The end-to-end multi-task perception framework was first trained and validated on the training and validation subsets of FaultData. Due to the limited strawberry growing season, the experiments were carried out indoors with artificial strawberries rather than in a strawberry orchard. During this evaluation, the relative error of the strawberry and the gripper beneath the picking point were measured, along with the physical error before compensation, the relative error before compensation, and the physical error after compensation, to assess picking accuracy. Furthermore, the end-to-end multi-task perception method was fine-tuned on SnapData-again initialized with the best weights from FaultData. The outputs of the multi-task perception model were then used as inputs to an LSTM classifier.
@@ 4.3 Results of Detection, Segmentation, Ripeness Estimation
SRR-Net, which integrated multi-task learning for both strawberry and hand classes, was evaluated on the FaultData dataset for detection, segmentation, and ripeness estimation in Table. 2. For bounding box prediction, it achieved a precision of 0.895, recall of 0.813, mAP@50 of 0.884, and mAP@50-95 of 0.633 for strawberries, while for the hand class, the corresponding values were 0.972, 0.958, 0.977, and 0.788, respectively. In segmentation, SRR-Net reached a precision of 0.887, recall of 0.747, mAP@50 of 0.829, and mAP@50-95 of 0.448 for strawberries, and 0.974, 0.947, 0.964, and 0.655 for hands.
@@
SRR-Net strawberry 0.999 0.992 0.995 0.978 0.999 0.992 0.995 0.981 - -
background 0.881 0.922 0.941 0.807 0.888 0.93 0.946 0.824 -
To more intuitively observe the performance of SRR-Net, the visualization results of SRR-Net, YOLOv11, and YOLOv11-seg on the FaultData dataset are presented in Fig. 10. In Fig. 10 (a) and (b), strawberries and end-effectors were represented with bounding boxes, with the target class label and confidence score displayed above each box. The confidence threshold was set to 0.5. For the same strawberry instance, SRR-Net produced higher confidence scores than YOLOv11 and YOLOv11-seg. In terms of ripeness estimation, ripeness estimation of SRR-Net closely matched the true ripeness of each strawberry. Overall, the visual comparisons illustrated that SRR-Net delivered superior performance compared with YOLOv11 and YOLOv11-seg.
@@ 4.4 Evaluation of Relative Error Compensation
From Table 3, the visual and ground-truth errors before compensation were compared, and the ground-truth errors after compensation were recorded. For example, in the first row, the relative errors on the $x$- and $y$- axis were 22 mm and -4 mm, while the corresponding ground-truth errors were 17.3 mm and 4.2 mm. The absolute differences between visual and ground-truth errors were therefore 4.7 mm and 8.2 mm. After compensation, the physical errors of the $x$- and $y$- axis were reduced to 1.5 mm and 6.3 mm, both below the defined threshold.
Before compensation, the mean relative errors were 14.07 mm on the $x$-axis and 8.64 mm on the $y$-axis, while the mean physical errors were 11.52 mm and 5.15 mm, respectively—indicating that visual estimation tended to overestimate the actual physical errors. With the proposed compensation method, the mean physical error between the strawberry point and the gripper were further reduced to 3.12 mm ($x$-axis) and 4.11 mm ($y$-axis). The slightly larger residual $y$-axis error was likely due to the difficulty of achieving high-precision motor control over very small movement ranges. Despite this phenomenon, the relative error compensation method demonstrated a clear advantage in aligning the gripper with the target and in reducing grasping errors.
Table 3: Relative errors between the end-effector and the picking point in the arm coordinate system (unit: $m$). The final row showed the computed mean absolute error.
@@ 4.5 Operational Efficiency Analysis of the Early Abort Strategy
To evaluate the effectiveness of the early abort strategy, the original picking time, the minimum time at which the early abort signal was triggered, and the time reduction for each action were calculated and analyzed in Table 4. In the original process, a complete picking cycle required approximately 12 s, with the inflating and approaching, swallowing, deflating, snap-off, descending, placing and homing stages taking 2 s, 2 s, 2 s, 1 s, 2 s, 2 s, and 1 s, respectively. When the gripper approached beneath the strawberry picking point, the relative error was estimated, and if it exceeded the threshold, a compensation action was triggered to align the gripper with the strawberry, adding 1 s to the cycle, as shown in Table. 4 #Case 1.
During the deflating stage, MobileNet V3-Small was executed to monitor whether a strawberry was present in the gripper. In accordance with the time-stability rule, the system required three consecutive signals before issuing the abort command. Upon receiving the abort command, the robotic arm immediately aborted the current process and skipped the descending and homing stages. The shortest time to receive the abort signal during the deflating stage was 0.11 s, resulting in a time reduction of 1.89 s, as shown in Table. 4 #Case 2. Overall, although the compensation action slightly increased the cycle time, the continuous harvesting efficiency was significantly improved.
For the snap-off stage, the results of the LSTM classifier on the validation set of SnapData were presented in Table. 5, where class 0 and class 1 represented strawberries remaining in the gripper and strawberries that had slipped, respectively. As shown in Table. 5, the LSTM classifier achieved high accuracy on SnapData. For class 0, all strawberries in the gripper were correctly detected (recall = 1.0000), with a few false positives (precision = 0.9294). For class 1, all predicted slips were correct (precision = 1.0000), though a small fraction were missed (recall = 0.9118). These results indicated that the LSTM classifier provided reliable slip prediction with minimal false alarms, supporting its use in early abort strategies. Classes 0 and 1 were classified using a maximum threshold. It was noted that class 0 encompassed two strawberry states—stable and slipping—so an adaptive minimum threshold was introduced and applied to further distinguish between them.
During the snap-off stage, operations were triggered based on the strawberry status in the gripper, with the time-stability rule applied to prevent false positives. When the early abort condition was met for three consecutive checks, the descending and placing stages were skipped, and homing was executed immediately. In randomized tests, the minimum response time for receiving an abort signal was 0.7 s, reducing the overall cycle time by 0.3 s, as presented in Table. 4 #Case 3.
Table 4: Harvesting time and reduction time via early abort of robot arm
Action Original Time (s) #Case 1 (s) #Case 2 (s) #Case 3 (s)
Relative Error Compensation Grasping failure Slipped
Inflation and approaching 2 2 2 2
Compensation - 1 - -
Swallowing 2 2 2 2
Deflating 2 2 0.11 2
Snap-off 1 1 - 0.7
Descending 2 2 2 2
Placing 2 2 - -
Homing 1 1 1 1
Time Reduction (s) - -1 1.89 0.3
Table 5: LSTM classifier results on the validation set of SnapData, where label 0 indicated that the strawberry remained in the gripper and label 1 indicated that it had slipped.
Class precision recall F1-score
0 0.9294 1.0000 0.9634
1 1.0000 0.9118 0.9538
Visualization results of the LSTM classifier on the validation of SnapData were presented in Fig. 11. The left panel displays frames 8-16 with the prediction probability and corresponding strawberry statuses, while the right panel illustrated the LSTM prediction probabilities. The horizontal axis represented the frame index starting from 0, and the vertical axis indicated the predicted probability of strawberry slippage within the next three frames. The red dashed line marked the maximum threshold, while the blue curve with asterisks showed the predicted probabilities. Red five-pointed stars denoted instances where the strawberry had slipped from the end-effector. An adaptive minimum threshold, illustrated by a pink dashed line, was employed to determine whether a secondary picking attempt should be executed or the subsequent actions should continue.
If the total number of frames did not reach 10, fewer than 10 frames were displayed as labels in the top-left corner of the images. The strawberry began slipping between frames 9 and 11, with corresponding probabilities of 0.406, 0.446, and 0.502. From frame 13 onward, the strawberry was considered slipped. Similarly, a probability curve of strawberry slippage was plotted for intuitive observation, as shown in the right panel of Fig. 11.
In Fig. 11 (left), the LSTM classifier presented the predicted probability of strawberry slippage for frames 8-16. When the number of frames did not reach 10, fewer than 10 frames were displayed as labels at the top-left of the images. The strawberry was slipping from frames 9-11, with corresponding probabilities of 0.406, 0.446, and 0.502; starting from frame 13, the strawberry had slipped. Similarly, a probability curve of strawberry slippage was plotted for intuitive observation in Fig. 11 right. The five-pointed star indicates the frame where the strawberry slipped for the first time. Based on validation results on SnapData, the LSTM classifier was able to timely send early-abort signals, enabling efficient and stable operation. Tests of the early abort strategy for empty grasp in deflating and slip prediction in snap-off were conducted on HarvestFlex. However, experiments involving inflating-and-deflating to re-grasp were not performed or deployed on HarvestFlex due to limitations in the strawberry growing season.
Figure 11: LSTM classifier results. (Left) Video frames 8-16; (Right) Predicted probabilities from the LSTM model.

### v2

@@ 3.4 Early Aborting Robotic Picking upon Grasp Adjustment and Slip Prediction
Empty grasp and misgrasp represented common failure modes in robotic strawberry harvesting, particularly during the deflating and snap-off stages. During the deflating stage, empty grasps occurred when the strawberry was not successfully enclosed by the end-effector, primarily due to localization errors or unexpected fruit motion. In the snap-off stage, strawberry slippage occurred when the picked strawberry unintentionally slipped, often due to a non-optimal end-effector pose or inadequate gripping force. To diagnose such faults effectively, a simple yet efficient method involved the installation of a miniature camera (JTS302, JiuTan, China) with a 3 mm diameter at the bottom of the end-effector, as shown in Fig. 7. The micro-optical camera directly captured real-time images of the gripper area, enabling visual monitoring of strawberry presence.
@@
Once the deflating stage began, a MobileNet V3-Small was used to determine whether a strawberry was present in the gripper, as shown in Fig. 8. If the classification result is 0 (strawberry correctly grasped), the robotic arm continued executing the subsequent picking actions. If the result is 1 (empty grasp) or 2 (misgrasp of unripe strawberry), the manipulator re-inflated the gripper, executed a descending motion, skipped the snapping and placing stages, and directly performed the reset motion to terminate the current picking cycle. Instead, it immediately selected a new target strawberry and proceeded to the next picking cycle. The process flow, indicated by the blue line, was illustrated in Fig. 4.
@@
For the slippage scenario, a time-series LSTM classifier was developed to predict whether the strawberry would slip from the end-effector during the snap-off stage. The LSTM classifier architecture was summarized in Table 1. The input layer of LSTM classifier had 7 features based a finetuned SRR-Net wtih strawberry and background classes. The LSTM layer consisted of 5 layers, a hidden size of 64, and a dropout of 0.2. This was followed by a Dropout layer with a rate of 0.3, then a Linear layer mapping from 64 to 3 units, and a Softmax layer applied along dimension 1. The output layer produced 3 classes.
@@
input 7
LSTM layers=5
hidden=64
dropout=0.2
Dropout 0.3
Linear (64, 3)
Softmax dim=1
output 3
The input of the LSTM classifier consisted of a sequence of features from the past 5 frames, represented as <<normalized strawberry area, normalized gripper area, normalized background area, $w$, $h$, $x$, $y>$. The normalized strawberry area, gripper area, and background area referred to the proportion of pixels in the image corresponding to the strawberry, the gripper, and the background, respectively. $w$ and $h$ represented the width and height of the strawberry. The model outputted the probability of slippage over the following three frames. If the probability fell below the minimum threshold, the strawberry remained attached to the end-effector after the snap-off motion, indicating a successful pick. Conversely, if the probability was equal to or exceeded the maximum threshold, the strawberry had slipped, indicating a failure in the snap-off stage. When the probability fell between the minimum and maximum thresholds, the strawberry was in the process of slipping from the gripper.
@@
To enhance reliability, a time-stability rule was implemented. If slipped was predicted in two consecutive instances, an early abort signal was triggered, the placing stage was skipped, and the next picking cycle began. If two consecutive frames were slipping, the end-effector retracted, reinflated, and re-performed the snap-off action. Otherwise, the robotic arm performed the subsequent actions.
@@ 3.5 Paradigm of HarvestFlex Visual Fault Diagnosis and Self-Recovery
The relative positional error compensation method, incorporating visual feedback SRR-Net, and the early abort strategy based on MobileNet V3-Small and the LSTM-based strawberry slip prediction were implemented as follows. The gripper then inflated and approached a location directly beneath the target strawberry picking point. At this position, the relative error between the gripper and the actual strawberry position was calculated. If the error exceeded a predefined tolerance threshold, a compensated picking point was generated to correct the deviation before proceeding.
Once aligned, the gripper executed swallowing and deflating motions to enclose the strawberry. During this process, a micro-optical camera mounted on the gripper captured images, which were used to determine whether the strawberry was successfully gripped based on MobileNet V3-Small classifier. When no strawberry or an unripe strawberry was detected in the gripper for two consecutive frames, the robot triggered re-inflation, descent, and homing actions to begin a new picking cycle. Conversely, it continuously performed the snap-off action while running a time-series LSTM classifier to predict strawberry normal.
To improve prediction stability, a time-stability rule was applied. When two consecutive slipped signals were triggered, the robot aborted the snap-off stage, skipped the placing stage, and returned to its homing position, as illustrated by the red lines in Fig. 10. When two consecutive normal signals were triggered, the robot continued the snap-off motion, as illustrated by the black line. Meanwhile, if two consecutive slipping signals were triggered, the gripper was inflated and deflated to re-grasp the strawberry during the deflating and snap-off actions, followed by the descending, placing, and homing steps, as illustrated by the orange line.
@@ 4.1 Evaluation Metrics
For the relative error between the end-effector and the picked strawberry instance, the visual relative errors along the $x$-, $y$-, and $z$- axes under the robotic arm coordinate system were denoted as $\Delta x$, $\Delta y$, and $\Delta z$, respectively. Due to the bending motion required during the snap-off stage, the end-effector’s ascent height could exceed that of the target point. Therefore, the error calculation was simplified from the $x$-, $y$-, and $z$-axes to only the $x$-, and $y$- axes. Similarly, the corresponding physical errors along the $x$-, and $y$-axes were denoted as $\Delta x_{w}$ and $\Delta y_{w}$. The physical error after compensation, denoted as $E_{x}$ and $E_{y}$, was measured to validate the effectiveness of the relative error compensation method. Finally, mean value $\bar{e}$ was used to evaluate the accuracy of the relative error and physical error before compensation and for each coordinate axis and can be formulated as:
@@
In addition, FPS (frames per second) was used to evaluate the inference speed of the model.
@@ 4.2 Implementation Details
To evaluate the effectiveness of the proposed method, experiments were conducted on the FaultData, GraspData, SnapData, and SlipData benchmarks. The end-to-end multi-task perception framework was first trained and validated on the training and validation subsets of FaultData. During this evaluation, the relative error of the strawberry and the gripper beneath the picking point were measured, along with the physical error before compensation, the relative error before compensation, and the physical error after compensation, to assess picking accuracy. Furthermore, the end-to-end multi-task perception method was fine-tuned using the images from the mini-camera. The outputs of the multi-task perception model were then applied to generate the SlipData for LSTM classifier. The model was trained using an NVIDIA RTX 4090 GPU.
@@ 4.3 Results of Detection, Segmentation, Ripeness Estimation
SRR-Net, which integrated multi-task learning for both strawberry and hand classes, was evaluated on the FaultData dataset for detection, segmentation, and ripeness estimation in Table. 2. For bounding box prediction, SRR-Net achieved a precision of 0.895, recall of 0.813, mAP@50 of 0.884, and mAP@50-95 of 0.633 for strawberries, while for the hand class, the corresponding values were 0.972, 0.958, 0.977, and 0.788, respectively. In segmentation, SRR-Net reached a precision of 0.887, recall of 0.747, mAP@50 of 0.829, and mAP@50-95 of 0.448 for strawberries, and 0.974, 0.947, 0.964, and 0.655 for hands.
@@
SRR-Net strawberry 0.991 0.99 0.992 0.98 0.991 0.99 0.994 0.984 - -
background 0.928 0.911 0.968 0.863 0.928 0.911 0.963 0.859 -
To more intuitively observe the performance of SRR-Net, the visualization results of SRR-Net, YOLOv11, and YOLOv11-seg on the FaultData dataset are presented in Fig. 11. In Fig. 11 (a) and (b), strawberries and end-effectors were represented with bounding boxes, with the target class label and confidence score displayed above each box. The confidence threshold was set to 0.5. The ripeness value of SRR-Net was displayed between the target class label and the confidence score. For the same strawberry instance, SRR-Net produced higher confidence scores than YOLOv11 and YOLOv11-seg. In terms of ripeness estimation, ripeness estimation of SRR-Net closely matched the true ripeness of each strawberry. Overall, the visual comparisons illustrated that SRR-Net delivered superior performance compared with YOLOv11 and YOLOv11-seg.
@@ 4.4 Evaluation of Relative Error Compensation
From Table 3, the visual and ground-truth errors before compensation were compared, and the ground-truth errors after compensation were recorded. For example, in the first row, the visual relative errors on the $x$- and $y$- axis were 22 mm and -4 mm, while the corresponding ground-truth errors were 17.3 mm and 4.2 mm. The absolute differences between visual and ground-truth errors were therefore 4.7 mm and 8.2 mm. After compensation, the physical errors of the $x$- and $y$- axis were reduced to 1.5 mm and 6.3 mm, both below the defined threshold.
Before compensation, the mean relative errors were 14.07 mm on the $x$-axis and 8.64 mm on the $y$-axis, while the mean physical errors were 11.52 mm and 5.15 mm, respectively—indicating that visual estimation tended to overestimate the actual physical errors. With the proposed compensation method, the mean physical error between the strawberry point and the gripper were further reduced to 3.12 mm ($x$-axis) and 4.11 mm ($y$-axis). The slightly larger residual $y$-axis error was likely due to the difficulty in achieving high-precision motor control over very small movement ranges. Despite this phenomenon, the relative error compensation method demonstrated a clear advantage in aligning the gripper with the target and in reducing grasping errors.
Table 3: Relative errors between the end-effector and the picking point in the arm coordinate system (unit: $m$). $(x_{s},y_{s},z_{s})$ denoted the coordinates of the initial picking point, and $(x_{e},y_{e},z_{e})$ represented the end-effector position when it moved to a location beneath the strawberry. The visual relative error between the vision-based picking point and the end-effector was defined as $\Delta x$ and $\Delta y$, while $\Delta x_{w}$ and $\Delta y_{w}$ denoted the corresponding positional error in the physical world. After applying the compensation strategy, the updated picking point was represented as $(x_{ce},y_{ce},z_{ce})$. The relative error between the compensated picking point and the end-effector in the physical world was denoted as $E_{x}$ and $E_{y}$. The final row showed the computed mean absolute error.
@@
4.5 Evaluation of Empty Grasp or Misgrasp
MobileNet-V3 Small was trained and validated based on the GraspData. The classification results were presented in Table 4. Class 0, 1, and 2 represent the presence of strawberries, the without strawberry, and the unripe strawberry in the end-effector, respectively. From the Table 4, it can be observed that the precision, recall, and F1-score for all classes are all 1.00, indicating that the model achieves optimal classification performance for each category and closely matches the ground-truth labels. The classification results for all categories were 1, mainly because the distinguishing features between missed grasps and incorrect grasps are relatively prominent. Overall, the classification results were excellent. To observe the classification results of MobileNet-V3 Small, the classification outputs of some images were visualized in Fig. 12. The upper-left corner of each image was annotated with the predicted class and the corresponding probability. From Fig. 12, it can be concluded that MobileNet-V3 Small achieved good classification performance.
Table 4: Results of empty grasp or misgrasp
Classes Precision Recall F1-score
0 1.00 1.00 1.00
1 1.00 1.00 1.00
2 1.00 1.00 1.00
Based on the classification results of MobileNet-V3 Small, different responses were triggered on the robotic arm. When the classification result was 0, the robotic arm continued to execute the subsequent operations. When the classification result was 1 or 2, the end-effector performed secondary inflation to prevent damage to the strawberry, then skipped the snap-off and placing stages and directly moved the end-effector downward to a position below the strawberry, thereby ending the current picking cycle.
(a) Strawberry in the end-effector
(b) Strawberry without in the end-effector
(c) Misgrasp of unripe strawberry
Figure 12: Visualization of empty grasp or misgrasp
4.6 Evaluation of Slip prediciton
Before performing LSTM classification in the snap-off stage, SRR-Net was first fine-tuned based on SnapData. These images contained two classes—strawberry and background—which were used to count the number of pixels belonging to the strawberry and background regions during the snap-off stage. The results of SRR-Net on SnapData were presented in Table 2. The precision and recall of the segmentation task reached 0.991 and 0.990 for strawberries, and 0.928 and 0.911 for the background, respectively, indicating that the model achieves highly accurate segmentation performance.
Table 5: LSTM classifier result on the SlipData
Class Precision Recall F1-socre
0 0.94 0.89 0.92
1 0.86 0.94 0.90
2 0.99 0.94 0.96
The normalized strawberry area, normalized background area, normalized hand area, width and height of strawberry, center point of strawberry as input of LSTM classifier to train and validate using SlipData. The results of the LSTM classifier on the validation set of SlipData were presented in Table. 5, where class 0, 1 and 2 represented strawberries remaining in the gripper, slipping and slipped, respectively. As shown in Table. 5, the LSTM classifier achieved high accuracy despite the imbalance among class samples.
Visualization results of the LSTM classifier on robots in the orchard were presented in Fig. 13. The left panel displayed frames 3-14 with the prediction probability, current strawberry status, the prediction probability of three status, and final strawberry status based on time-stability, while the right panel illustrated the LSTM prediction probabilities. The horizontal axis represented the frame index starting from 0, and the vertical axis indicated the predicted probability of strawberry slippage within the next three frames. If the total number of frames did not reach 5, fewer than 10 frames were displayed as labels in the top-left corner of the images. The LSTM classifier predicted a slipping status at frames 7, 11, and 12, with corresponding probabilities of 0.62, 0.70, 0.82. Based on the time-stability criterion, secondary inflating and snap-off were triggered at frame 12. The probability curve of strawberry slip prediction was plotted for intuitive observation, as shown in the right panel of Fig. 13. Similarly, a probability curve of strawberry slippage was plotted for intuitive observation in Fig. 13 right.
Figure 13: LSTM classifier results. (Left) Video frames 3-14; (Right) Predicted probabilities from the LSTM model. Note that frames were counted starting from 0.
@@ 4.7 Operational Efficiency Analysis of the Early Abort Strategy
[…]
Table 6: Harvesting time of robot arm (unit: s). – indicated that the execution time at the corresponding stage under a certain fault was identical to the original (baseline) time. $\times$ indicated that the corresponding stage was not executed.
Action Original Empty Grasp Misgrasp Slipped Compensation
Inflating & approaching 1.25 $\pm$ 0.01
Compensation $\times$ $\times$ $\times$ $\times$ 0.71 $\pm$ 0.07
Swallowing 0.73 $\pm$ 0.00
[…]
Descending 0.98 $\pm$ 0.08
Placing 4.36 $\pm$ 0.01 $\times$ $\times$ $\times$
Homing 1.88 $\pm$ 0.00
[…]
For grasp adjustment, empty grasp times and misgrasp time were shown in Fig. 15. The response times for empty grasp and misgrasp were 0.42 $\pm$ 0.04 and 0.39 $\pm$ 0.03, respectively. Compared with original deflating stage, grasp adjustment can save approximately 0.5 s. The snap-off and placing stages were skipped, which saved time in these two phases. Compared with baseline method, the mitigation of grasp adjustment reduced the harvesting cycle time by approximately 4 seconds, thereby improving overall harvesting efficiency in Table. 6.
Figure 14: Relative Error Compensation Time
Figure 15: Response time of grasp adjustment during the deflating stage (unit: s).
For strawberry slip prediction, the harvesting time for strawberries with normal status was 1.03 $\pm$ 0.00 s. For strawberry slipping that triggered re-inflating and re-snap off, the execution time was 1.81 $\pm$ 0.07 s. For slipped, the total time was 1.44 $\pm$ 0.07 s. It should be noted that the time recorded for the confirmed slippage state includes the duration of re-inflation and re-bending attempts, followed by the final detachment of the strawberry. During the snap-off stage, if the strawberry was detected to be slipping, the completion time for secondary inflation, grasping signal generation, and manipulator response was 1.81 $\pm$ 0.07 s. When the strawberry had already slipped off, the manipulator response time decreased to 1.44 $\pm$ 0.07 s. With early termination enabled, the total execution time of the harvesting cycle was reduced to 7.08 $\pm$ 0.10 s. Compared with the original cycle time of 11.21 $\pm$ 0.15 s, the proposed strategy reduced the harvesting cycle time by approximately 4 s. The picking cycle time increased to 12.51 $\pm$ 0.09 s after integrating relative error compensation for grasping and slip prediction, compared with the original cycle time of 11.21 $\pm$ 0.15 s, representing an increase of approximately 1.3 s.
Table 7: Robot arm response time of strawberry slip prediction (unit: s).
Strawberry Status Normal Slipping Slipped
Time 1.03 $\pm$ 0.00 1.81 $\pm$ 0.07 1.44 $\pm$0.07
Furthermore, the success rates of grasp verification and strawberry slippage prediction were assessed through repeated trials based on the robotic arm’s response. As shown in Table. 8, the system achieves an approximately 90.00% success rate in identifying empty grasp or grasp of unripe strawberry, thereby ensuring appropriate corrective actions on the robot arm. For strawberry slip prediction during the snap-off stage, the success rate for strawberries in a normal state reached 96.88%. For slipping, the success rate of secondary grasping was 81.25%, while slipped was 88.89%. The relatively lower success rate for initiating secondary inflation and snap-off during snap-off was mainly attributed to the short duration of the slip event, which requires rapid prediction within a limited time window. In addition, the performance was influenced by the initial grasping posture of the end-effector during the first contact with the strawberry.
Table 8: Statistics on robotic arm response success rate for repeated grasp judgments and slip predictions
Result Grasp Adjustment Strawberry slip prediction
Empty Grasp Misgrasp Normal Slipping Slipped
Success 27 30 31 26 32
Failure 3 3 1 6 4
Success Rate 90.00% 90.91% 96.88% 81.25% 88.89%